# APEX2: Adaptive and Extreme Summarization for Personalized Knowledge Graphs

### v1

@@ A.1. GLIMPSE
1:  Compute $\mathcal{T}^{\Delta\neq 0}$ with Pr$(e|\mathcal{Q})$, Pr$(x_{ijk}|\mathcal{Q})$
2:  $\mathcal{P}\leftarrow\emptyset$
@@ A.2. PEGASUS
1:  $\mathcal{V}_{P}\leftarrow\{\{u\}:u\in\mathcal{V}\}$; $\mathcal{E}_{P}\leftarrow\{\{\{u\},\{v\}\}:\{u,v\}\in\mathcal{E}\}$
2:  $t\leftarrow 1;\theta\leftarrow 0.5;\mathcal{L}\leftarrow[]$
@@ A.4. APEX2
1:  $\mathbf{H}\leftarrow|\mathcal{E}|\times|\mathcal{R}|\times|\mathcal{E}|$ Sparse Matrix with 0s, $\mathcal{E}_{p}^{(0)}\leftarrow\emptyset$, $\mathcal{R}_{p}^{(0)}\leftarrow\emptyset$, $\mathcal{T}_{p}^{(0)}\leftarrow\emptyset$
2:  for $i=0$; $i<d;i\texttt{++}$+ do
@@
7:  $\mathcal{E}_{p}^{(0)}\leftarrow\{e\in\mathcal{E}\quad{\rm s.t.}\quad\exists x\in\mathcal{T}_{p}^{(0)},e\in x\}$ and $\mathcal{R}_{p}^{(0)}\leftarrow\{r\in\mathcal{R}\quad{\rm s.t.}\quad\exists x\in\mathcal{T}_{p}^{(0)},r\in x\}$ /* Adapting Phase */
8:  for timestamp $t=1$; $t\leq T$; $t\texttt{++}$+ do
@@ A.5. APEX2-N
1:  $\mathbf{H}\leftarrow|\mathcal{E}|\times|\mathcal{E}|$ Sparse Matrix with 0s, $\mathcal{E}_{p}^{(0)}\leftarrow\emptyset$, $\mathcal{R}_{p}^{(0)}\leftarrow\emptyset$, $\mathcal{T}_{p}^{(0)}\leftarrow\emptyset$
@@
7:  $\mathcal{T}_{p}^{(0)}\leftarrow\{x_{ijk}\in\mathcal{T}\quad{\rm s.t.}\quad i,j\in\mathcal{E}_{p}^{(0)},v\in e\}$
8:  $\mathcal{R}_{p}^{(0)}\leftarrow\{r\in\mathcal{R}\quad{\rm s.t.}\quad\exists x_{ijk}\in\mathcal{T}_{p}^{(0)},r\in x_{ijk}\}$ /* Updating Phase */
9:  for timestamp $t=1$; $t\leq T$; $t\texttt{++}$+ do
@@
11:     $\mathcal{C}$ = HeatDiffuse$(\mathbf{H},\mathcal{G},\mathcal{Q}^{(t)}\setminus\mathcal{Q}^{(t-1)},d)$
@@ B.1.2. PEGASUS
(25) $\begin{split}RE^{(\mathcal{T})}(\mathcal{P})=\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}\bm{W}_{ij}^{(\mathcal{T})}|\bm{A}_{ij}^{(\mathcal{G})}-\bm{A}_{ij}^{(\hat{\mathcal{G}})}|\end{split}$
where $\hat{\mathcal{G}}=(\mathcal{V},\hat{\mathcal{E}})$ is the reconstructed graph from $\mathcal{P}$. $\bm{A}_{ij}$ stands for the adjacency matrix. $\mathcal{T}$ is the given target node set as interests. $\bm{W}_{ij}^{\mathcal{(T)}}$ is the personalized weight of each pair of nodes $(i,j)$. $\bm{W}_{ij}^{\mathcal{(T)}}$ depends on the pairwise distance to the target nodes,
@@ C.1. Metric: F1 Score
[…]
@@
[…]
@@ C.4.1. Code
The code for the main experiment is provided at
https://drive.google.com/file/d/1rgQgtrLmcLRbcPL1h8EIPa3boKhdkz5p/view?usp=sharing. The data of DBpedia and MetaQA is provided within our submission code, but YAGO3 is not due to its large size. You need to manually download YAGO3 follow our instruction if you want to play with it. Our code only supports ”.gz” Compressed Archive Folders, so you need to convert using gzip if the downloaded files are in other format.
@@ C.6. Experiments Adjusting $R_{APEX}$
APEX2-$1$ 0.631 0.055
APEX2-$2$ 0.586 0.049
APEX2-$3$ 0.540 0.044
APEX2-$6$ 0.426 0.036
APEX2-N-$1$ 0.858 0.018
APEX2-N-$2$ 0.849 0.011
APEX2-N-$3$ 0.766 0.007
APEX2-N-$6$ 0.680 0.004
@@
APEX2-$1$ 0.601 6.354
APEX2-$2$ 0.562 4.947
APEX2-$3$ 0.524 3.433
APEX2-$6$ 0.475 2.695
APEX2-N-$1$ 0.882 2.528
APEX2-N-$2$ 0.861 1.691
APEX2-N-$3$ 0.772 1.052
APEX2-N-$6$ 0.667 0.689
@@ Appendix E Proof of Theorems
(33) ($\begin{split}{{k+n-1}\choose{n-1}}\cdot k!=\frac{(k+n-1)!}{(n-1)!\cdot k!}\cdot k!=\frac{(n+k-1)!}{(n-1)!}\end{split}$
@@
Respectively, for each timestamp $\mathbf{e}[i],\mathbf{r}[j]$ and $\mathbf{e}[k]$ has on average $2$, $1$, and $\frac{2nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l}))}{|\mathcal{E}|}$ entries updated, therefore $\mathbf{H}$ has on average $O(3\cdot 2\cdot 2\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{2})=O(12\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{2})$ entries changed. The incremental sorting algorithm on $\mathbf{H}$ will take $O(12\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{2}\cdot\log_{2}nnz(\mathbf{H}))<O(12\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{2}\cdot\log_{2}(12\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{3}))=O(36\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{2}\cdot\log_{2}(12\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|))$. ∎
@@
(37) $\begin{split}\sum_{x_{ijk}\in\mathcal{T}_{u}}\log{\rm Pr}(x_{ijk}|\mathcal{Q})&<\sum_{x_{ijk}\in\mathcal{T}_{v}}\log{\rm Pr}(x_{ijk}|\mathcal{Q})\\ \prod_{x_{ijk}\in\mathcal{T}_{u}}{\rm Pr}(e_{i}|\mathcal{Q}){\rm Pr}(r_{k}|\mathcal{Q}){\rm Pr}(e_{j}|\mathcal{Q})&<\prod_{x_{ijk}\in\mathcal{T}_{v}}{\rm Pr}(e_{i}|\mathcal{Q}){\rm Pr}(r_{k}|\mathcal{Q}){\rm Pr}(e_{j}|\mathcal{Q})\\ \mathbb{E}_{e,u}^{|\mathcal{E}_{u}|+2|\mathcal{T}_{u}|}\mathbb{E}_{r,u}^{|\mathcal{T}_{u}|}&<\mathbb{E}_{e,v}^{|\mathcal{E}_{v}|+2|\mathcal{T}_{v}|}\mathbb{E}_{r,v}^{|\mathcal{T}_{v}|}\\ \end{split}$
@@
(41) $\begin{split}\sum_{e\in\mathcal{E}_{u}}\log{\rm Pr}(e|\mathcal{Q})&<\sum_{e\in\mathcal{E}_{v}}\log{\rm Pr}(e|\mathcal{Q})\\ \prod_{e\in\mathcal{E}_{u}}{\rm Pr}(e|\mathcal{Q})&<\prod_{e\in\mathcal{E}_{v}}{\rm Pr}(e|\mathcal{Q})\\ \mathbb{E}_{e,u}^{|\mathcal{E}_{u}|}&<\mathbb{E}_{e,v}^{|\mathcal{E}_{v}|}\\ (\frac{\gamma^{b}-\gamma^{a+b}}{1-\gamma}\cdot 2\cdot\frac{1-(\alpha c_{u})^{d+1}}{1-\alpha c_{u}}\cdot\frac{1}{|\mathcal{E}_{u}|})^{|\mathcal{E}_{u}|}&<(\frac{1-\gamma^{b}}{1-\gamma}\cdot 2\cdot\frac{1-(\alpha c_{v})^{d+1}}{1-\alpha c_{v}}\cdot\frac{1}{|\mathcal{E}_{v}|})^{|\mathcal{E}_{v}|}\\ \end{split}$

### v2

@@ A.1. GLIMPSE
1: Compute $\mathcal{T}^{\Delta\neq 0}$ with $(e|\mathcal{Q}), Pr$$(x_{ijk}|\mathcal{Q})$
[…]
@@ A.2. PEGASUS
[…]
@@ A.4. APEX2
[…]
2:for $i=0$; $i<d;i\texttt{++}$ do
@@
[…]
8:for timestamp $t=1$; $\leq T; t++t\texttt{++}$ do
@@ A.5. APEX2-N
[…]
@@
[…]
9:for timestamp $t=1$; $\leq T; t++t\texttt{++}$ do
@@
[…]
@@ B.1.2. PEGASUS
[…]
@@ C.1. Metric: F1 Score
(28) $\textit{TP(True Positive)}=|\mathcal{A}_{pkg}\cap\mathcal{A}_{kg}|$
(29) $\textit{FP(False Positive)}=|\mathcal{A}_{pkg}\setminus\mathcal{A}_{kg}|$
(30) $\textit{FN(False Negative)}=|\mathcal{A}_{kg}\setminus\mathcal{A}_{pkg}|$
@@
(32) $\textit{F1}=\frac{2\times\textit{precision}\times\textit{recall}}{\textit{precision}+\textit{recall}}$
@@ C.4.1. Code
The code for the main experiment is provided at https://github.com/iDEA-iSAIL-Lab-UIUC/APEX. The data of DBpedia and MetaQA is provided within our submission code, but YAGO3 is not due to its large size. You need to manually download YAGO3 follow our instruction if you want to play with it. Our code only supports ”.gz” Compressed Archive Folders, so you need to convert using gzip if the downloaded files are in other format.
@@ C.6. Experiments Adjusting $R_{APEX}$
APEX2-11 0.631 0.055
APEX2-22 0.586 0.049
APEX2-33 0.540 0.044
APEX2-66 0.426 0.036
APEX2-N-11 0.858 0.018
APEX2-N-22 0.849 0.011
APEX2-N-33 0.766 0.007
APEX2-N-66 0.680 0.004
@@
APEX2-11 0.601 6.354
APEX2-22 0.562 4.947
APEX2-33 0.524 3.433
APEX2-66 0.475 2.695
APEX2-N-11 0.882 2.528
APEX2-N-22 0.861 1.691
APEX2-N-33 0.772 1.052
APEX2-N-66 0.667 0.689
@@ Appendix E Proof of Theorems
(33) $\begin{split}{{k+n-1}\choose{n-1}}\cdot k!=\frac{(k+n-1)!}{(n-1)!\cdot k!}\cdot k!=\frac{(n+k-1)!}{(n-1)!}\end{split}$
@@
Respectively, for each timestamp $\mathbf{e}[i],\mathbf{r}[j]$ and $\mathbf{e}[k]$ has on average $2$, 11, and $\frac{2nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l}))}{|\mathcal{E}|}$ entries updated, therefore $\mathbf{H}$ has on average $O(3\cdot 2\cdot 2\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{2})=O(12\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{2})$ entries changed. The incremental sorting algorithm on $\mathbf{H}$ will take $O(12\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{2}\cdot\log_{2}nnz(\mathbf{H}))<O(12\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{2}\cdot\log_{2}(12\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{3}))=O(36\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|^{2}\cdot\log_{2}(12\cdot\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}\cdot|\mathcal{Q}|))$. ∎
@@
[…]
@@
[…]